OpenStudy (popstar30977):

how do you divide 175/35 (show an example)

Vocaloid (vocaloid):

(without using a calculator, long division involves a bit of guessing and checking) first, we have to estimate how many times 35 goes into 175 without going over I'll start by guessing 4 35*4 = 140, which is less than 175 so we need to go up now I'll guess 5 35*5 = 175 so 175/35 = 5

OpenStudy (popstar30977):

@Vocaloid I did not get it

Vocaloid (vocaloid):

what are you still confused about?

OpenStudy (popstar30977):

wait i got it never mind :) thanks for the help

OpenStudy (popstar30977):

@Vocaloid